OpenStudy (anonymous):

segment with endpoints A (4, 2) and C (1, 5) is partitioned by a point B such that AB and BC form a 1:3 ratio. Find B

OpenStudy (anonymous):

(1, 2.5) (2.5, 3.5) (3.25, 2.75) (3.75, 4.5)

OpenStudy (anonymous):

@ganeshie8

ganeshie8 (ganeshie8):

use the section formula, do u have the formula in ur notes ?

OpenStudy (anonymous):

no

ganeshie8 (ganeshie8):

Point dividing a given segment in \(l : m \) ratio :- \(\huge \left( \frac{l x_2 + m x1}{l+m} , \frac{l y_2 + m y_1}{l+m} \right)\)

ganeshie8 (ganeshie8):

plugin the values

ganeshie8 (ganeshie8):

A (4, 2) and C (1, 5) x1 y1 x2 y2

OpenStudy (anonymous):

whats l and m

ganeshie8 (ganeshie8):

\(l\) = 1 m = 3